Sally bought three chocolate bars and a pack of gum and paid $1.75. Jake bought two chocolate bars and four packs of gum and paid $2.00. Find the cost of a chocolate bar and the cost of a pack of gum.
Answer by checkley71(8403)

3C+G=1.75 OR G=1.75-3C
2C+4G=2
2C+4(1.75-3C)=2
2C+7-12C=2
-10C=2-7
-10C=-5
C=-5/-10
C=1/2 OR $.50 ANSWER.
G=1.75-3*.50
G=1.75-1.50
G=.25 ANSWER.
PROOF
2*.50+4*.25=2
1+1=2
2=2